1.You wish to order penicillin VK for little Jerry who weighs 35 lbs. The recommended dosage is 56 mg/kg/day divided into 4 doses for 10 days. You plan to use Pen VK 125mg/5ml. How much should he receive for each dosage?

2. Paul is an 11-month-old who you want to prescribe a corticosteroid. His weight is 14 pounds, 6 ounces. You will prescribe Pediapred 6.7 mg/5ml. You wish to give 2 mg/kg/day in three divided dosages for 5 days. What will you order?

3. Andrew is a 2-year-old who you want to prescribe amoxicillin. He weighs 12.5 kilograms pounds. You will use amoxicillin 250mg/5 ml. You want to give 50mg/kg/day in three divided doses for 10 days.

4. Polly weighs 45 pounds. You want to give her amoxicillin/clavulanate (Augmentin). The recommended dosage is 20 mg/kg/day of amoxicillin in three divided doses for ten days. You will use 200mg amoxicillin/28.5mg clavulanate/ 5 ml. Calculate the dosage based on the amoxicillin component.

5. Karen weighs 29 pounds. You want to order Zithromax, 10mg/kg/d for day 1 then 5 mg/kg/day on days 2 through 5. You will use 200mg/5ml. What will you order?

6. James is a 4-year-old who weighs 35 kilograms. You want to prescribe erythromycin for him. You will use erythromycin 400 mg/5ml. The dosage is 50mg/kg/day in 4 divided dosages for 14 days. What will you order?

7. Mia is a 5-year-old who weighs 65 pounds. You want to order clarithromycin 250 mg/5 ml for her. The dosage is 15 mg/kg/day in two divided dosages for 10 days. What will you order?

8. Jerry is newly retired and complaining of insomnia. He is 66-years-old and weighs 210 pounds. What medication and dose would you prescribe for him.
Answer #1

1. order dose penicillin VK  56 mg/kg/day divided into 4 doses for 10 days.

Jerry weighs 35 lbs (15.8kg).

You plan to use Pen VK 125mg/5ml

means 25mg/ml

The dose will be

56mg/kg\times 15.8kg=884.8mg

Jerry will receive

884.8mg\div 25mg/ml=35.3ml/day

If you divide it into 4 similar doses then you have to give 8,8ml/dose

ANSWER- Jerry will receive 8.8ml/dose

2.

Order dose 2 mg/kg/day in three divided dosages for 5 days.

His weight is 14 pounds, 6 ounces (6.52kg)

You will prescribe Pediapred 6.7 mg/5ml.

The total dose will be

2mg/kg\times 6.52kg=13.04 mg/day

You will give it in 3 divide dosage, then you have to give 4.34 mg/dose

Available drug is 6.7mg/5ml (1.3mg/ml)

6.7mg\div 5ml = 1.3mg/ml

So,

4.34mg\div 1.3mg/ml= 3.3ml/dose

Answer - Nurse will order 3.3ml /dose

3.

Order dose - 50mg/kg/day in three divided doses for 10 days.

He weighs 12.5 kilograms.

You will use amoxicillin 250mg/5 ml.

(50mg/ml)

He will receive

50mg/kg\times 12.5kg=625mg/day

If you give 3 doses per day then he will receive 208.3mg/dose

So, he will receive

208.3mg\div 50mg/ml=4.1ml/dose

Answer - He will receive 4.1ml/dose

4.

Polly weighs 45 pounds(20.4 kg).

The recommended dosage is 20 mg/kg/day of amoxicillin in three divided doses for ten days.

You will use 200mg amoxicillin/28.5mg clavulanate/ 5 ml.

Polly will receive

20mg/kg\times 20.4kg = 408mg

Then

200mg amoxicillin in 5 ml. (40mg/ml)

408mg\div40mg/ml = 10.2ml

Answer - Polly will receive 10.2ml